OpenStudy (anonymous):

how do you find the distance between two points using thee Pythagorean theorem?

Vocaloid (vocaloid):

draw a right triangle

Vocaloid (vocaloid):



Vocaloid (vocaloid):



Vocaloid (vocaloid):

then use the Pythagorean theorem to find c (which is the distance between the two points) a^2 + b^2 = c^2

OpenStudy (hyroko):

\[x ^{2}+y ^{2}= distance ^{2}\]

OpenStudy (anonymous):

thanks guys so much, and any one who else wants to answer it feel free lol

OpenStudy (hyroko):

the distance between your X's squared plus the distance between your Y's squared will give you the distance between the points squared.

OpenStudy (hyroko):

(x2-x1)^2 +(y2-y1)^2 = distance^2

OpenStudy (anonymous):

@Hyroko so for x2 and y2 could that be a2 nd y2?

OpenStudy (hyroko):

First point (x1,y1) Second point (x2,y2)

OpenStudy (hyroko):

so x2-x1 will give you the distance between the points on the X axis and y2-y1 will give you the distance between the points on the Y axis. then you square those values. then add them together. then take the square root of your answer and you have the distance between the 2 points.

OpenStudy (hyroko):

for example, if you had a point (0,5) and the point (5,6) 5-0 = 5 6-5 = 1 (5)^2 + (1)^2 = 26 the distance would be \[\sqrt{26}\]

OpenStudy (hyroko):

just realized that may be a bad example.... Lol (3,4) (5,6) 5-3 = 2 6-4 = 2 (2)^2+(2)^2 = 8 distance is \[\sqrt{8}\]

OpenStudy (anonymous):

thanks so much @Hyroko